OpenStudy (anonymous):

Consider both Earth and Mars to be spheres. The surface area of mars is about 28% that of earth. What is the ratio of their radii?

OpenStudy (asnaseer):

do you know the formula for the surface area of a sphere?

OpenStudy (anonymous):

4pir^2

OpenStudy (asnaseer):

correct, so lets call the radius of Mars \(r_m\) and the radius of Earth \(r_e\)

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

you are given:\[\frac{4\pi r_m^2}{4\pi r_e^2}=\frac{28}{100}\]

OpenStudy (asnaseer):

and you need to work out what \(\displaystyle\frac{r_m}{r_e}\) is from this.

OpenStudy (asnaseer):

can you do it now?

OpenStudy (anonymous):

if 4pir^2 is the area of a sphere then the surface area of mars compared to that of earth is 0.28*(4pir^2) if earth's is just 4pir^2 right. equate both sides and solve for r. Asnaseer's answer might be more thorough but if you know how to solve for r here with basic algebra that's all there is to it.

OpenStudy (anonymous):

yeah I forgot how you solve for r

OpenStudy (anonymous):